OpenStudy (tennistar):

Multiply: 2 the square root of 11 • 8 the square root of 22

OpenStudy (tennistar):

I got 3 questions wrong so I have to retake my test

OpenStudy (anonymous):

\[2\sqrt{11}*8\sqrt{22}\] this one?

OpenStudy (tennistar):

yes

OpenStudy (anonymous):

show me where you stuck

OpenStudy (tennistar):

well I get to\[16\sqrt{242}\] and I get stuck

OpenStudy (anonymous):

that's ok, just need to simpify the expretion under the root.

OpenStudy (tennistar):



OpenStudy (tennistar):

answer choices 64 the square root of 22 10 the square root of 33 8 the square root of 11 176 the square root of 2

OpenStudy (anonymous):

From your simplifying, you see that to get 121 you just multiply 11*11, so the square root of 121 is 11. So: \[16 \sqrt{242}=16\sqrt{121} \sqrt{2} = 16*11\sqrt{2}\] Finish that and you have the answer.

OpenStudy (anonymous):